Here is the problem: I don't really know for what type of shape the surface area is for...
The small intestine has a length of 22 feet and a 2-inch circumference. Find its surface area in square feet.

Answer by the_tutor2011(8)

shape is a cylinder

C = 2pi*r, C = 2
2 = 2pi*r
pi*r = 1
r = 1/pi
r = 0.3183098861838

SA = 2pi*r^2 + 2pi*r*h

SA = 2(3.14)(0.3183098861838)^2 + 2pi*(0.3183098861838)*(12*22)

SA = 528.63629703329

Surface area is 528.63629703329 square ft, multiply by 144 to convert to sq in.


528.63629703329*144 = 76,123.6267727938


Answer is 76,123.6267727938 sq in